OpenStudy (anonymous):

Find the rate of change for 2000 to 2010 given these ordered pairs in the data set: (2000, 286) and (2010, 192). Write your answer as a decimal. A. -0.94 B. -9.4 C. -94 D. -940

OpenStudy (anonymous):

What's the formula for RATE of change?

OpenStudy (anonymous):

@Joshuam1 respond to me!

OpenStudy (anonymous):

Use the two points to find the slope.

OpenStudy (anonymous):

slope = (192 - 286)/(2010 - 2000) = - 94/10 = is what

OpenStudy (anonymous):

Sorry @Chlorophyll I walked out. But that's my problem, I can't remember the formulas

OpenStudy (anonymous):

Thanks, Zoe.